OpenStudy (anonymous):

Please learn me how to do this 2x+3y=8 x-3y=-5 What does x equal?

OpenStudy (anonymous):

make Y the subject of formula in both of the equations then equate the two equations and find the value of x

OpenStudy (anonymous):

I do not understand what you mean.

OpenStudy (texaschic101):

It might be better to isolate x in the 2nd equation and substitute

OpenStudy (anonymous):

please dumb it down for me

OpenStudy (anonymous):



OpenStudy (texaschic101):

I take that back....elimination is probably easier in this case

OpenStudy (texaschic101):

2x + 3y = 8 x - 3y = -5 ------------add 3x = 3 -- divide both sides by 3 x = 1

OpenStudy (anonymous):

There you go two methods to do it :3 i hope you understand

OpenStudy (anonymous):

thanks much

OpenStudy (texaschic101):

thats what we are here for :)

OpenStudy (uri):

2x+3y=8 x-3y=-5 ======== 3x=3 x=3/3 x=1 +3y and -3y gets cancelled.

OpenStudy (uri):

@Jack_RM

OpenStudy (anonymous):

I see